# OmniLens++: Blind Lens Aberration Correction via Large LensLib Pre-Training and Latent PSF Representation

### v1

@@ 3.2 Data Preparation Stage: Large LensLib Construction
Hybrid sampling basis. Since traditional design indicators do not fully capture the optical degradation of a lens in the image domain (§ 3.1), we quantify optical degradation by assessing the image quality of a lens’s imaging results. We quantify per‑FoV severity and categorize optical degradation by its variation trend to enable balanced sampling covering possible degradation patterns. Concretely, as in Figure 2 (b), a degraded checkerboard of the target lens and its paired GT are applied as the quantification base. Five knife-edge image patches sampled from the center to periphery FoVs are cropped for quantification. The Optical Image Quality (OIQ) is assessed to grade per‑FoV degradation severity, where we adopt the weighted sum of fidelity metrics (PSNR and SSIM) and optical metric (SFR) as a prototype of OIQ. The average OIQ across the $5$ FoVs is applied as the overall degradation severity of the target lens. For the entire lens source, we analyze the per‑sample average OIQ distribution and partition the severity into $3$ classes as Strong, Medium, and Mild. Furthermore, we categorize spatial variation patterns of optical degradation by defining $6$ Optical Degradation Class (OD‑Class) from OIQ trends over FoVs. First, using the variance and mean of per‑FoV OIQ, we compute the coefficient of variation to measure spatial uniformity $U_{S}$. Samples with values above a threshold $\alpha$ are labeled the “spatial-uniform” OD‑Class. For $U_{S}<\alpha$, $5$ additional OD-Class are defined by the FoV of peak OIQ and the monotonicity of OIQ changes. In this way, average OIQ and OD-Class serve as the hybrid sampling basis for sampling a uniform OD distribution. Detailed OIQ computation and OD-Class definitions are provided in Appendix D.
@@ 4.1 Implementation Details
Datasets. We sample $m_{1}=200$ and $m_{2}=3$ instances per class for AODLibpro train and AODLibpro test, yielding $3,600$ training lenses and $54$ test lenses with no overlap. For AODLibpro train, each lens randomly degrades $40$ images from Flickr2K (timofte2017ntire) via a precise imaging simulator (yang2024end). For AODLibpro test, the degraded images are synthesized based on additional $26$ collected clear images. Meanwhile, through manually designing and gathering open source designs, we construct a real-world lens aberration dataset RealLens-Sim, consisting of $4$ minimalist optical systems with spherical or aspheric surfaces (MOS-A/S) (jiang2024flexible), $1$ metalens from tseng2021neural(MOS-Meta), one smartphone lens under $3$ misalignment (MA) cases, and $2$ high-end lenses including an ultra-wide lens and a smartphone lens (High-end). The same simulator is applied to simulate the paired images of them for numerical evaluation. For real-world cases, we collect real-snapped optical degradation images RealLens-Snap with fabricated single lenses, the commercial minimalist lens CAYECAYE $50mm$ $f1/4$, two DSLRs of CanonCanon $24mm$ $f1/4$ and SonySony $18135$, and nano-optics data from tseng2021neural.
Training details. On AODLibpro train, we first pre-train LPR for $100K$ iterations, and train the FoundCAC model for $200K$ iterations. Both training stages use $256\times 256$ random crops with flips and rotations for data augmentation, with a batch size of $16$. We use Adam with a learning rate of 2e42e-4 to 1e61e-6 decayed by a cosine annealing schedule. Training is conducted on two RTX 4090 GPUs, and inference uses a single RTX 4090. Consistent with chen2022real, all encoders and decoders employ $3$ groups of ResBlocks. For FoundCAC, we use $4$ RSTB layers from Swin-T (liang2021swinir) as the deep fusion backbone, given its strong performance on spatial-varying OD (jiang2024minimalist). Ablations on alternative backbones are reported in Section 4.2 and 4.4.
@@ Appendix C Detailed Settings for Experiments in Motivation
The schematic diagrams of the model paradigms in Table 1 of Section 3.1 are shown in Figure 6. All encoders and decoders use the same design as in the final FoundCAC, namely $3$ groups of ResBlocks. All fusions between PSF features and image features adopt $F_{CAC}$, that is, concatenation followed by deep feature backbones of $4$ RSTB layers. Supervision for both PSFs prediction and PSFs features prediction applies L1 loss. Other training settings follow those of the CAC stage in FoundCAC. Additionally, all models in this table are trained on AODLibpro-Train and evaluated on the AODLibpro-Test benchmark to evaluate network performance.
@@ D.2 Details for hybrid sampling basis
where PSNR and SSIM are processed following (liang2024ntire) to obtain normalized metrics, while SFR is represented as OIQE (jiang2024minimalist), which denotes the ratio of the SFR of the evaluated target to that of a lens without OD. The weights are set to $\lambda_{1}=0.4,\lambda_{2}=0.3,\lambda_{3}=0.3$ to control the value ranges so that the $3$ normalized metrics are close in range following (liang2024ntire).
@@
Definition of OD severity category. The average OIQ across the $5$ knife-edge image patches of different FoVs is calculated to represent the overall severity level of the target lens’s OD. The average OIQ lies within $[0,1]$, where a larger value indicates lower overall OD severity (higher optical image quality). Therefore, we divide the range into $3$ intervals to categorize different OD severity levels for sampling, as shown in Figure 7.
Calculation of the spatial uniformity. We first define the coefficient of variation (CV) from the variance and mean of OIQ across the $5$ FoVs, then use it to compute a spatial uniformity metric $U_{S}$ that measures the uniformity of the OD spatial distribution:
@@
Discussion on chromatic aberrations. We do not include chromatic aberration as a criterion in OD classification because preliminary experiments show that it has little impact on the final CAC results, as shown in Figure 9. Using the same computation pipeline as $U_{S}$, we compute the channel-wise uniformity of OIQ to quantify the severity of chromatic aberration and divide it into $5$ categories, where a smaller category index indicates more severe chromatic aberration. The distribution of chromatic aberration in AODLibpro Test is shown in Figure 9 (a). We tally the performance of our trained FoundCAC at different chromatic levels and find no obvious pattern in its performance as the severity of chromatic aberration changes, as shown in Figure 9 (b). This indicates that chromatic aberration has little effect on CAC models trained on LensLib, even though it is an often-discussed OD pattern. Meanwhile, our experiments also include cases with pronounced chromatic aberration, as in Figure 4, where our method effectively suppresses purple fringing caused by chromatic aberration. These evidences indicate that although chromatic aberration is not considered in data sampling, it does not affect the model’s performance in this respect.
@@ E.1 Details for PSF-VQVAE
The encoder and decoder of PSF-VQVAE each use $3$ groups of ResBlocks, with corresponding downsampling or upsampling convolutions in every group. The latent space for representation and feature processing operates at $1/8$ of the input resolution and has $n_{z}=256$ channels. The structures of the encoder and decoder and the latent space feature dimensionality apply equally to ODN and FoundCAC, so we do not elaborate further. The codebook contains $K=1024$ code vectors with a feature dimension of $512$. During feature matching, two additional $1\times 1$ convolution layers are applied before and after matching to handle feature dimension mismatches. These designs and settings follow chen2022real; esser2021taming.
@@ E.3 Details for CAC stage
For FoundCAC (RRDB), we use $6$ RRDB layers from wang2018esrgan as the deep fusion backbone. For FoundCAC (Swin), we adopt $4$ RSTB layers in liang2021swinir. For the RRDB layer, num_grow_chnnum\_grow\_chn is set to $128$, corresponding to a feature channel dimension of $256$. For the RSTB layer, we set blk_depthblk\_depth, num_headsnum\_heads, and window_sizewindow\_size to $6$, 88, and $8$, respectively. Since our RSTB operates at $1/8$ input resolution, FoundCAC achieves faster inference than SwinIR, which uses RSTB at $1/4$ resolution. In $F_{CAC}$, we adopt concatenation as the first step because our goal is to demonstrate that LPR can effectively guide CAC, and concatenation followed by deep fusion is an intuitive approach. Exploring more efficient fusion strategies is also an interesting direction for future work.
@@ F.1 Details on imaging simulation
For AODLibpro Train, AODLibpro Test, and RealLens-Sim, we obtain the paired OD images corresponding to each lens design via imaging simulation. The imaging simulator from DeepLens (yang2024end) is adopted, considering its precise calculation of PSFs and simulation of OD images by patch-wise convolution. Specifically, we feed the design files of each sample in these LensLibs (in Zemax or parameter table format) into DeepLens for ray tracing to compute PSFs at $64$ fields of view and $31$ sampled wavelengths in the visible band. We match the closest sensor from the sensor library ($4\mu{m}-2K$, $8\mu{m}-2K$, $12\mu{m}-2K$, $16\mu{m}-2K$) based on the image height, discretize and sample the PSFs according to the pixel size, and stack them across RGB channels according to the wavelength response characteristics, ultimately obtaining the PSF arrays for each FoV and channel used in the simulation. It is worth noting that we applied equivalent downsampling to sensor resolution and pixel size to facilitate the use of large‑scale public $2K$ high‑quality image datasets. Meanwhile, unifying the resolution also helps to systematically control variables to build a benchmark for exploring the CAC task. Finally, the computed PSFs are used to convert clear images into the corresponding OD images via patch-wise convolution, while the pipeline also accounts for sensor ISP and noise, as in most imaging simulation workflows (chen2021extreme_quality; chen2021optical). For each lens in the $3$ datasets, we will open source its Zemax design file, the computed PSF array, and the simulated paired OD images.
@@ F.3 Illustration of lens designs in RealLens-Sim
Figure 11 shows the structures of the test lenses used in RealLens‑Sim and example imaging results. The lenses are sourced from open source designs in the literature or from designs manually created by optical designers based on specifications in public patents or the needs of minimalist applications. In addition, we consider lens misalignment, a common but rarely addressed real-world factor that induces OD. This typically occurs when the PSF size under the nominal lens design is below one pixel, that is, no OD, but the decentering and tilt errors in manufacturing and assembly lead to random unknown OD in the final imaging results. We select a smartphone lens whose original design yields no sensor sampled OD, set $3$ groups of random decentering and tilt errors with increasing magnitude within its tolerance range, and then perform ray tracing to compute PSFs for imaging simulation.
@@ F.4 Illustration of capture details for RealLens-Snap
We show the imaging setups used for our real-world captures in RealLens-Snap in Figure 12. In addition, the nano-optics data come from $3$ open-source images in tseng2021neural, and the CanonCanon $4mm$ $f1/4$ data come from two open-source images in eboli2022fast. Quantitative evaluation on RealLens-Snap directly reflects the potential of blind aberration correction to improve image quality on real-world terminals. Given the difficulty of collecting real-world lenses under various applications for shooting, we strive to construct RealLens-Snap covering commonly used scenarios for blind aberration correction, such as minimalist optical systems, high-end photographic equipment, and metalens imaging. The selected systems exhibit distinct OD distributions, enabling a more comprehensive evaluation of the representative methods. To the best of our knowledge, we are the first work of blind aberration correction whose evaluation simultaneously covers minimalist optical systems with severe aberrations and high‑end lenses with mild aberrations. We also hope to continuously collect more real snapped images with OD in future work to broaden application scenarios.

### v3

@@ 3.2 Data Preparation Stage: Large LensLib Construction
Hybrid sampling basis. Since traditional design indicators do not fully capture the optical degradation of a lens in the image domain (§ 3.1), we quantify optical degradation by assessing the image quality of a lens’s imaging results. We quantify per‑FoV severity and categorize optical degradation by its variation trend to enable balanced sampling covering possible degradation patterns. Concretely, as in Figure 2 (b), a degraded checkerboard of the target lens and its paired GT are applied as the quantification base. Five knife-edge image patches sampled from the center to periphery FoVs are cropped for quantification. The Optical Image Quality (OIQ) is assessed to grade per‑FoV degradation severity, where we adopt the weighted sum of fidelity metrics (PSNR and SSIM) and optical metric (SFR) as a prototype of OIQ. The average OIQ across the 55 FoVs is applied as the overall degradation severity of the target lens. For the entire lens source, we analyze the per‑sample average OIQ distribution and partition the severity into 33 classes as Strong, Medium, and Mild. Furthermore, we categorize spatial variation patterns of optical degradation by defining 66 Optical Degradation Class (OD‑Class) from OIQ trends over FoVs. First, using the variance and mean of per‑FoV OIQ, we compute the coefficient of variation to measure spatial uniformity $U_{S}$. Samples with values above a threshold $\alpha$ are labeled the “spatial-uniform” OD‑Class. For $U_{S}<\alpha$, 55 additional OD-Class are defined by the FoV of peak OIQ and the monotonicity of OIQ changes. In this way, average OIQ and OD-Class serve as the hybrid sampling basis for sampling a uniform OD distribution. Detailed OIQ computation and OD-Class definitions are provided in Appendix D.
@@ 4.1 Implementation Details
Datasets. We sample $m_{1}=200$ and $m_{2}=3$ instances per class for AODLibpro train and AODLibpro test, yielding $3,600$ training lenses and $54$ test lenses with no overlap. For AODLibpro train, each lens randomly degrades $40$ images from Flickr2K (timofte2017ntire) via a precise imaging simulator (yang2024end). For AODLibpro test, the degraded images are synthesized based on additional $26$ collected clear images. Meanwhile, through manually designing and gathering open source designs, we construct a real-world lens aberration dataset RealLens-Sim, consisting of 44 minimalist optical systems with spherical or aspheric surfaces (MOS-A/S) (jiang2024flexible), 11 metalens from tseng2021neural(MOS-Meta), one smartphone lens under 33 misalignment (MA) cases, and 22 high-end lenses including an ultra-wide lens and a smartphone lens (High-end). The same simulator is applied to simulate the paired images of them for numerical evaluation. For real-world cases, we collect real-snapped optical degradation images RealLens-Snap with fabricated single lenses, the commercial minimalist lens CAYECAYE $50mm$ $f1/4$, two DSLRs of CanonCanon $24mm$ $f1/4$ and SonySony $18135$, and nano-optics data from tseng2021neural.
Training details. On AODLibpro train, we first pre-train LPR for $100K$ iterations, and train the FoundCAC model for $200K$ iterations. Both training stages use $256\times 256$ random crops with flips and rotations for data augmentation, with a batch size of $16$. We use Adam with a learning rate of 2e42e-4 to 1e61e-6 decayed by a cosine annealing schedule. Training is conducted on two RTX 4090 GPUs, and inference uses a single RTX 4090. Consistent with chen2022real, all encoders and decoders employ 33 groups of ResBlocks. For FoundCAC, we use 44 RSTB layers from Swin-T (liang2021swinir) as the deep fusion backbone, given its strong performance on spatial-varying OD (jiang2024minimalist). Ablations on alternative backbones are reported in Section 4.2 and $4.$
@@ Appendix C Detailed Settings for Experiments in Motivation
The schematic diagrams of the model paradigms in Table 1 of Section 3.1 are shown in Figure 6. All encoders and decoders use the same design as in the final FoundCAC, namely 33 groups of ResBlocks. All fusions between PSF features and image features adopt $F_{CAC}$, that is, concatenation followed by deep feature backbones of 44 RSTB layers. Supervision for both PSFs prediction and PSFs features prediction applies L1 loss. Other training settings follow those of the CAC stage in FoundCAC. Additionally, all models in this table are trained on AODLibpro-Train and evaluated on the AODLibpro-Test benchmark to evaluate network performance.
@@ D.2 Details for hybrid sampling basis
where PSNR and SSIM are processed following (liang2024ntire) to obtain normalized metrics, while SFR is represented as OIQE (jiang2024minimalist), which denotes the ratio of the SFR of the evaluated target to that of a lens without OD. The weights are set to $\lambda_{1}=0.4,\lambda_{2}=0.3,\lambda_{3}=0.3$ to control the value ranges so that the 33 normalized metrics are close in range following (liang2024ntire).
@@
Definition of OD severity category. The average OIQ across the 55 knife-edge image patches of different FoVs is calculated to represent the overall severity level of the target lens’s OD. The average OIQ lies within $[0,1]$, where a larger value indicates lower overall OD severity (higher optical image quality). Therefore, we divide the range into 33 intervals to categorize different OD severity levels for sampling, as shown in Figure 7.
Calculation of the spatial uniformity. We first define the coefficient of variation (CV) from the variance and mean of OIQ across the 55 FoVs, then use it to compute a spatial uniformity metric $U_{S}$ that measures the uniformity of the OD spatial distribution:
@@
Discussion on chromatic aberrations. We do not include chromatic aberration as a criterion in OD classification because preliminary experiments show that it has little impact on the final CAC results, as shown in Figure 9. Using the same computation pipeline as $U_{S}$, we compute the channel-wise uniformity of OIQ to quantify the severity of chromatic aberration and divide it into 55 categories, where a smaller category index indicates more severe chromatic aberration. The distribution of chromatic aberration in AODLibpro Test is shown in Figure 9 (a). We tally the performance of our trained FoundCAC at different chromatic levels and find no obvious pattern in its performance as the severity of chromatic aberration changes, as shown in Figure 9 (b). This indicates that chromatic aberration has little effect on CAC models trained on LensLib, even though it is an often-discussed OD pattern. Meanwhile, our experiments also include cases with pronounced chromatic aberration, as in Figure 4, where our method effectively suppresses purple fringing caused by chromatic aberration. These evidences indicate that although chromatic aberration is not considered in data sampling, it does not affect the model’s performance in this respect.
@@ E.1 Details for PSF-VQVAE
The encoder and decoder of PSF-VQVAE each use 33 groups of ResBlocks, with corresponding downsampling or upsampling convolutions in every group. The latent space for representation and feature processing operates at $1/8$ of the input resolution and has $n_{z}=256$ channels. The structures of the encoder and decoder and the latent space feature dimensionality apply equally to ODN and FoundCAC, so we do not elaborate further. The codebook contains $K=1024$ code vectors with a feature dimension of $512$. During feature matching, two additional $1\times 1$ convolution layers are applied before and after matching to handle feature dimension mismatches. These designs and settings follow chen2022real; esser2021taming.
@@ E.3 Details for CAC stage
For FoundCAC (RRDB), we use 66 RRDB layers from wang2018esrgan as the deep fusion backbone. For FoundCAC (Swin), we adopt 44 RSTB layers in liang2021swinir. For the RRDB layer, num_grow_chnnum\_grow\_chn is set to $128$, corresponding to a feature channel dimension of $256$. For the RSTB layer, we set blk_depthblk\_depth, num_headsnum\_heads, and window_sizewindow\_size to 66, 88, and 88, respectively. Since our RSTB operates at $1/8$ input resolution, FoundCAC achieves faster inference than SwinIR, which uses RSTB at $1/4$ resolution. In $F_{CAC}$, we adopt concatenation as the first step because our goal is to demonstrate that LPR can effectively guide CAC, and concatenation followed by deep fusion is an intuitive approach. Exploring more efficient fusion strategies is also an interesting direction for future work.
@@ F.1 Details on imaging simulation
For AODLibpro Train, AODLibpro Test, and RealLens-Sim, we obtain the paired OD images corresponding to each lens design via imaging simulation. The imaging simulator from DeepLens (yang2024end) is adopted, considering its precise calculation of PSFs and simulation of OD images by patch-wise convolution. Specifically, we feed the design files of each sample in these LensLibs (in Zemax or parameter table format) into DeepLens for ray tracing to compute PSFs at $64$ fields of view and $31$ sampled wavelengths in the visible band. We match the closest sensor from the sensor library ($4\mu{m}-2K$, $8\mu{m}-2K$, $12\mu{m}-2K$, $16\mu{m}-2K$) based on the image height, discretize and sample the PSFs according to the pixel size, and stack them across RGB channels according to the wavelength response characteristics, ultimately obtaining the PSF arrays for each FoV and channel used in the simulation. It is worth noting that we applied equivalent downsampling to sensor resolution and pixel size to facilitate the use of large‑scale public $2K$ high‑quality image datasets. Meanwhile, unifying the resolution also helps to systematically control variables to build a benchmark for exploring the CAC task. Finally, the computed PSFs are used to convert clear images into the corresponding OD images via patch-wise convolution, while the pipeline also accounts for sensor ISP and noise, as in most imaging simulation workflows (chen2021extreme_quality; chen2021optical). For each lens in the 33 datasets, we will open source its Zemax design file, the computed PSF array, and the simulated paired OD images.
@@ F.3 Illustration of lens designs in RealLens-Sim
Figure 11 shows the structures of the test lenses used in RealLens‑Sim and example imaging results. The lenses are sourced from open source designs in the literature or from designs manually created by optical designers based on specifications in public patents or the needs of minimalist applications. In addition, we consider lens misalignment, a common but rarely addressed real-world factor that induces OD. This typically occurs when the PSF size under the nominal lens design is below one pixel, that is, no OD, but the decentering and tilt errors in manufacturing and assembly lead to random unknown OD in the final imaging results. We select a smartphone lens whose original design yields no sensor sampled OD, set 33 groups of random decentering and tilt errors with increasing magnitude within its tolerance range, and then perform ray tracing to compute PSFs for imaging simulation.
@@ F.4 Illustration of capture details for RealLens-Snap
We show the imaging setups used for our real-world captures in RealLens-Snap in Figure 12. In addition, the nano-optics data come from 33 open-source images in tseng2021neural, and the CanonCanon $4mm$ $f1/4$ data come from two open-source images in eboli2022fast. Quantitative evaluation on RealLens-Snap directly reflects the potential of blind aberration correction to improve image quality on real-world terminals. Given the difficulty of collecting real-world lenses under various applications for shooting, we strive to construct RealLens-Snap covering commonly used scenarios for blind aberration correction, such as minimalist optical systems, high-end photographic equipment, and metalens imaging. The selected systems exhibit distinct OD distributions, enabling a more comprehensive evaluation of the representative methods. To the best of our knowledge, we are the first work of blind aberration correction whose evaluation simultaneously covers minimalist optical systems with severe aberrations and high‑end lenses with mild aberrations. We also hope to continuously collect more real snapped images with OD in future work to broaden application scenarios.